OpenStudy (anonymous):

You are hiking along the California coast and wonder about the height of a particular Giant Redwood tree. You are 5 feet and nine inches tall and your shadow is 11 feet long. The shadow of the tree is 440 feet long. How tall is the tree?

OpenStudy (amistre64):

congruent triangles are congruent...or something like that. (23/4) : 11 as x:440 x = (23/4)(440)/11

OpenStudy (amistre64):

x = (23)(110)/(11) = (23)(10) = 230... if I did the math right

OpenStudy (anonymous):

THANKS !

OpenStudy (anonymous):

Yea that sounds right!

OpenStudy (anonymous):

where did you get 23 /4 ? just so i can understand it better